OpenStudy (anonymous):

Find the surface area of a sphere that has a volume of 288pi cu.in a)48pi b)144pi c)20,736pi

OpenStudy (sparrow2):

v=4/3pr^3 so in your case r=6(just enter the rumbers in the formula) than surface area is 4pr^2 so we have 4px36=144pi

OpenStudy (anonymous):

Oh! Thank you Sparrow2!

OpenStudy (triciaal):

welcome

OpenStudy (michele_laino):

hint: the volume of a sphere of radius R is: \[V = \frac{{4\pi }}{3}{R^3}\]

OpenStudy (michele_laino):

so we can write: \[{R^3} = \frac{{3V}}{{4\pi }}\]

OpenStudy (michele_laino):

now you have to solve that last equation for R, what do you get?

OpenStudy (triciaal):